What is the solution set of |3x + 4| <= 7
Answer by khwang(438)

Use |ax +b|<= c is equivalent to -c <= ax+b <= c
We see that |3x + 4| <= 7 is equivalent to
-7 <= 3x+4 <= 7
Subtract 4 on each term, we have
-11 <= 3x <= 3,
Divide each term by 3, we get
-11/3 <= x <= 1.
Hence, the solutin set is [-11/3, 1].